OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 4x – y = 4 x + 4y = 18

OpenStudy (anonymous):

Get one of the variables in one of the equations by itself. 4x - y = 4 -y = 4 - 4x y = 4x - 4 Now substitute (4x - 4) into the second equation for y, and solve for x. x + 4(4x - 4) = 18 x + 16x - 16 = 18 17x = 34 x = 2 Now substitute 2 for x in one of the equations and solve for y. 2 + 4y = 18 4y = 16 y = 4 Final answer: x = 2 y = 4 You are welcome. :)

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

No problem (And I'm a girl) lol :)

OpenStudy (anonymous):

:P sorry

OpenStudy (anonymous):

Haha, it's okay.